OpenStudy (kitkat77):

2 • {[5 • (60 – 14 ÷ 7)] + 25} • 5

jabez177 (jabez177):

Now this one is similar to the other one.

jabez177 (jabez177):

But it has Brackets.

OpenStudy (anonymous):

543.571428571

jabez177 (jabez177):

You do the Brackets first and whatever is inside the brackets.

OpenStudy (wendy.ivette11714):

This problem is similar, but it has brackets. This is how you solve it: =2(5(60−2)+25)(5) =2((5)(58)+25)(5) =2(290+25)(5) =(2)(315)(5) =(630)(5) =3150

OpenStudy (kitkat77):

Okay thank you all so much you have helped me so much.

OpenStudy (wendy.ivette11714):

My pleasure

OpenStudy (kitkat77):

:)